40m QIL Cryo_Lab CTN SUS_Lab CAML OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 7113     Entry time: Wed Aug 8 09:46:29 2012
Author: Masha 
Type: Update 
Category: Environment 
Subject: ETMX EQ 

[Sasha, Masha, Liz, Eric]


A bunch of surfs in the lab just noticed that ETMX is going crazy (laser is shifting everywhere) due to a 4.5 EQ that just hit LA. The optic is already shut down according to the watchdogs.

ELOG V3.1.3-